To demonstrate flavor aversion learning (that is, learning to dislike a flavor that is associated with becoming sick), researchers gave one group of laboratory rats an injection of lithium chloride immediately following consumption of saccharin-flavored water. Lithium chloride makes rats feel sick. A second control group was not made sick after drinking the flavored water. The next day, both groups were allowed to drink saccharin-flavored water. The amounts consumed (in milliliters) for both groups during this test are given below. Amount Consumed by Rats That Were Made Sick (n = 4) Amount Consumed by Control Rats (n = 4) 1 7 5 12 4 11 3 9 (a) Test whether or not consumption of saccharin-flavored water differed between groups using a 0.05 level of significance. State the value of the test statistic. (Round your answer to three decimal places.) Incorrect: Your answer is incorrect. State the decision to retain or reject the null hypothesis. Retain the null hypothesis. Reject the null hypothesis. Correct: Your answer is correct. (b) Compute effect size using eta-squared (η2). (Round your answer to two decimal places.) η2 = Incorrect: Your answer is incorrect.
Answer #1
Amount Consumed by Rats That Were Made Sick (n = 4) Amount Consumed by Control Rats (n = 4)
1 4
7 11
5 3
12 9
Mean 6.25 6.75
SD 4.573474245 3.862210075

Testing for Equality of Variances AF-test is used to test for the equality of variances. The following F-ratio is obtained: 4The provided sample means are shown below: X1 6.25 X2 =6.75 Also, the provided sample standard deviations are 814.57 82 = 3.8

(2) Rejection Region Based on the information provided, the significance level is a 0.05, and the degrees of freedom are df 6(5) Conclusion It is concluded that the null hypothesis Ho is not rejected. Therefore, there is not enough evidence to claim

Effect size = Mean diff / Sp

s2p = (SS1 + SS2) / (df1 + df1)

s2p = ((df1)(s21) + (df2)(s22)) / (df1 + df2) = 35.78 / 6 = 17.892

Sp =  4.2299

Effect size = (6.75-6.25)/ 4.2299 = 0.1182